# HyConEx: Hypernetwork classifier with counterfactual explanations

### v1

@@ 3.4 HyConEx architecture
To explain our approach, we assume that $x\in\mathbb{R}^{D}$ was classified by $f(x;W)$ as the element of the class $k$. It means that the $k$-th row of $W$ is related to the vector that is normal to the local decision boundary generated for $x$. To localize the counterfactual examples of alternative classes, we employ the remaining rows of $W$. More precisely, the translation of $x$ by the $m$-th row $W_{m}$, defined by $x^{\prime}=x-W_{m}$, has to point out the counterfactual example of the class $m$. To satisfy this requirement, we need to ensure that $x^{\prime}$:
@@
    is classified by $f$ as the element of $m$-th class,
@@
    is close to the original input $x$,
@@
    is plausible (in-distribution sample).
First, the correct class prediction of $x^{\prime}$ is forced by minimizing the cross-entropy loss, $\mathcal{L}(f(x^{\prime};W),m)$ using our hypernetwork classifier $W=H(x^{\prime};\theta)$, see Figure 2(b). Second, to satisfy the proximity between $x$ and $x^{\prime}$, we minimize the MSE between them, i.e. $\|x-x^{\prime}\|^{2}$, see Figure 2(c). Third, the plausibility is obtained using class conditional INFs. More precisely, we train class conditional INF, $F(x|y;\phi)$, to model the density of each class. Next, we adjust the counterfactual example $x^{\prime}$ to align with a fixed value of the estimated data density, see Figure 2(d). In our paper, we follow the PPCEF [38] approach and put:
$\mathcal{L}_{F}(x^{\prime},m)=\max\left(\delta-p_{F}(x^{\prime}|m),0\right),$
where $p_{F}(x^{\prime}|m)$ denotes the class conditional probability of the counterfactual explanation $x^{\prime}$ under the desired target class value $m$ and $\delta$ represents the density threshold.
The complete loss function of HyConEx combines the above three loss terms with the cross-entropy of the original input $x$ to provide the correct classification. Since we are interested in generating counterfactual examples for all alternative classes, we iterate over all classes. Let us first define the counterfactual loss for $m$-th class:
$\mathcal{L}_{conEx}(x,m;\theta,\phi)=\alpha_{1}\mathcal{L}(f(x^{\prime};W),m)+\alpha_{2}\|x-x^{\prime}\|+\alpha_{3}L_{F}(x^{\prime},m),$
where $x^{\prime}=x-W_{m}$ and $\alpha_{i}$’s are regularization hyperparameters. Then our final loss equals:
$\mathcal{L}_{HyConEx}(x,y;\theta,\phi)=\mathcal{L}(f(x),y)+\sum_{m\neq y}\mathcal{L}_{conEx}(x,m;\theta,\phi).$
@@ 3.5 Initialization and training process of HyConEx
Training of HyConEx involves training of hypernetwork $H(\cdot;\theta)$ together with conditional INF, $F(\cdot;\phi)$. For stability, we first pre-train both networks as described below.
Since the hypernetwork produces not only a classification boundary but also counterfactual explanations, we cannot pre-train the hypernetwork only with a typical cross-entropy loss. For the hypernetwork to be aware of the closest elements of opposite classes, we apply the clustering of each class separately using the k-means algorithm [8]. Then we combine the classification cross-entropy loss $\mathcal{L}(f(x),y)$ with the proximity loss between counterfactual candidate $x^{\prime}$ and the nearest clustering mean of the alternative class:
$\mathcal{L}(f(x),y)+\alpha\sum_{m\neq y}\|x^{\prime}-r_{m}\|$
where $x^{\prime}=W_{m}$ and $r_{k}$ is the closest cluster center from the $k$-th class to a data point $x$ and $\alpha$ is a hyperparameter. This clustering-based pre-training is also targeted to obtain plausible counterfactuals, but with lower quality than applying INFs. However, it is more stable which is important at the initial stage of the training.
We also train a conditional INF $F(\cdot;\phi)$ on class labels predicted by a pre-trained hypernetwork classifier. This initialization helps to stabilize the training process and leads to faster convergence.
After pre-training stage, we freeze the parameters of INF and train only the hypernetwork network until the convergence.
@@ 4 Experiments
We evaluate HyConEx in terms of classification accuracy and counterfactual explanation quality. To our knowledge, HyConEx is the first deep learning model that learns to return counterfactual examples by a classification model itself.
@@ 4.1 Instantiation of HyConEx
For implementing HyConEx, we use TabResNet as hypernetwork backbone and MAF as conditional INF. TabResNet uses 4 residual blocks, a hidden size of 256, and a dropout rate of 0.25. MAF is configured with 16 hidden features, 8 layers, and 4 blocks per layer.
The initialization phase uses trade-off parameter $\alpha=0.8$ and is continued until model convergence. In the second phase, we employ the same Adam optimizer and cosine annealing scheduler as used during pre-training. For all experiments, we linearly increase the trade-off parameters $\alpha_{i}$ from 0 to $\alpha_{1}$=0.8, $\alpha_{2}$=0.1 and $\alpha_{3}$=0.1. The duration of the second phase is determined using early stopping, which aims to maximize plausibility while ensuring high model accuracy, high counterfactual validity, and low proximity.
Table 1: Comparison of predictive performance using AUROC measure.
Dataset Decision Logistic Random TabNet TabResNet CatBoost IMN HyConEx
ID Tree Regression Forest
3 0.987 0.990 0.998 0.983 0.999 0.999 0.999 0.997
12 0.938 0.999 0.998 0.995 0.999 0.999 0.999 0.999
31 0.643 0.775 0.795 0.511 0.756 0.790 0.751 0.799
54 0.804 0.938 0.927 0.501 0.968 0.934 0.957 0.946
1067 0.620 0.802 0.801 0.789 0.808 0.800 0.805 0.809
1461 0.703 0.908 0.930 0.926 0.931 0.937 0.930 0.905
1464 0.599 0.749 0.666 0.516 0.740 0.709 0.742 0.769
1468 0.926 0.996 0.995 0.495 0.995 0.996 0.994 0.996
1486 0.935 0.987 0.993 0.991 0.994 0.994 0.993 0.987
1489 0.842 0.805 0.962 0.928 0.949 0.948 0.950 0.900
1590 0.752 0.903 0.917 0.908 0.915 0.930 0.915 0.872
40984 0.946 0.980 0.995 0.985 0.994 0.995 0.994 0.985
41142 0.626 0.742 0.796 0.713 0.782 0.822 0.775 0.725
41143 0.749 0.850 0.880 0.823 0.860 0.870 0.865 0.835
41146 0.910 0.966 0.983 0.974 0.982 0.988 0.981 0.979
41161 0.857 0.995 0.999 0.997 0.998 1.000 0.998 0.994
41163 0.873 0.994 0.999 0.998 1.000 1.000 1.000 0.992
41164 0.786 0.898 0.925 0.888 0.913 0.935 0.902 0.895
@@ 4.2 Predictive performance
To validate the predictive performance of HyConEx, we use the benchmark from [21]. More precisely, we compare HyConEx with strong ensemble methods (Random Forest and CatBoost), classical shallow classifiers (decision trees, logistic regression) as well as deep learning models equipped with interpretability tools (TabNet [4], TabResNet, IMN [2]). We use 18 tabular datasets retrieved from OpenML repository of varied characteristics (see Table 1 for datasets IDs). We report AUROC (Area Under the Receiver Operating Characteristic curve), which is well suited for imbalanced data as well.
The results presented in Table 1 demonstrate that HyConEx maintains competitive performance compared to state-of-the-art methods, achieving the highest AUROC on several datasets (31, 1067, 1464, 1468). In particular, HyConEx shows comparable performance to TabResNet and CatBoost, suggesting that the additional counterfactual generation capability does not compromise classification accuracy. This confirms that our model effectively preserves predictive performance while gaining interpretability through counterfactual explanations.
@@ 4.3 Counterfactual explanations
We evaluated counterfactual explanations by assessing the quality, plausibility, and computational efficiency of our HyConEx method compared to established baselines such as CEM [12], CBCE [23], CEGP [35], and PPCEF [38]. Our evaluation leverages datasets from the PPCEF benchmark [38] and expands the original PPCEF setup by incorporating additional categorical datasets. We consider multiple scenarios, including binary classification using datasets such as Moons, Law, Audit and Heloc, multiclass classification using datasets such as Blobs, Digits and Wine, and cases with categorical features using Adult [7], Credit-A [30] and Credit-G [18] datasets. For all experiments, we perform standard preprocessing: normalizing features and handling class imbalance via downsampling. For datasets with categorical features we introduce slight Gaussian noise to labels for efficient training, while for counterfactual examples, we apply softmax activation to produce one-hot predictions and select the class with maximum value.
@@
Metrics
We quantify counterfactual quality using several metrics. Coverage (Cover.) indicates the fraction of instances for which counterfactuals are generated, while Validity (Valid.) represents the percentage of counterfactuals that successfully changed the model prediction to the desired target. To assess the minimality of the changes from the original example, we report the L1 and L2 norms for continuous features and Hamming distance for categorical features. Plausibility is measured using a probabilistic plausibility score (P.Plaus.) which indicates how many counterfactuals have an estimated log density higher than the median log density of the train set. Furthermore, we provide density metrics such as log density (LogDens) and outlier detection scores such as Local Outlier Factor (LOF) and Isolation Forest (IsoForest). Finally, we report the CPU computation time in seconds to generate counterfactuals to capture the efficiency of the method.
@@
Results
Our experiments demonstrate that HyConEx consistently achieves near-perfect coverage and validity across diverse settings while offering substantial computational advantages. Results presented in Table 2 show that for binary classification tasks, our method generates counterfactual explanations with competitive proximity metrics and high plausibility, processing the moons dataset in merely 0.003 seconds compared to 0.75 seconds for CBCE. This pattern extends to multiclass scenarios as shown in Table 3 and datasets with categorical features as illustrated in Table 4, where HyConEx maintains strong performance while producing valid counterfactuals that require only minimal modifications and remain within high-density regions of the training data distribution. By providing counterfactual explanations that are both valid and plausible, our approach enhances the practical interpretability and actionability of explanations while significantly outperforming baseline methods in computation speed, demonstrating its effectiveness for counterfactual analysis across binary, multiclass, and categorical tabular data.

### v2

@@ 4.1 Architecture
To explain our approach, we assume that $x\in\mathbb{R}^{D}$ is classified by $f(x;W)$ as the element of the class $k$. It means that the $k$-th row of $W$ is related to the vector that is normal to the local decision boundary generated for $x$. To localize the counterfactual examples for any other class $m$, we utilize the remaining rows of $W$, denoted as $W_{m}$. Concretely, we define a counterfactual candidate $x^{\prime}$ as a translated version of x, given by $x^{\prime}=x-W_{m}$. For $x^{\prime}$ to indeed represent a valid counterfactual instance of class $m$, the following conditions must be met:
@@
    The classifier $f(\cdot;H(x^{\prime};\theta))$ assigns $x^{\prime}$ to class $m$,
@@
    $x^{\prime}$ is close to the original input $x$,
@@
    $x^{\prime}$ is plausible (in-distribution sample).
The first criterion is satisfied by minimizing the cross-entropy loss $\mathcal{L}_{CE}(f(x^{\prime};W^{\prime}),m)$, where $W^{\prime}=H(x^{\prime};\theta)$ is the weight matrix returned by the hypernetwork $H$ for instance $x^{\prime}$ (as shown in Figure 2(b)). Second, to satisfy the proximity between $x$ and $x^{\prime}$, we minimize the MSE between them, i.e. $\|x-x^{\prime}\|^{2}$, see Figure 2(c). And the final one is addressed using class-conditional invertible normalizing flow $F(\cdot|m;\phi)$, with parameters $\phi$, which estimates the data density within each class separately (as illustrated in Figure 2(d)). Following the approach proposed in PPCEF [10], we employ a cost function that encourages $x^{\prime}$ to reach a fixed density threshold. This function is given by:
$\mathcal{L}_{F}(x^{\prime},m;\phi)=\max\left(\delta-p_{F}(x^{\prime}|m;\phi),0\right),$ (6)
where $p_{F}(x^{\prime}|m)$ represents the probability density of the counterfactual example $x^{\prime}$, conditioned on class $m$, and $\delta$ is a threshold set based on the median density of the training data.
By combining these three cost components into a single expression, we obtain the counterfactual loss $\mathcal{L}_{conEx}$ for a given alternative class $m$:
$\mathcal{L}_{conEx}(x,x^{\prime},m;\theta,\phi)=\alpha_{1}\mathcal{L}_{CE}(f(x^{\prime};H(x^{\prime};\theta)),m)+\alpha_{2}\|x-x^{\prime}\|+\alpha_{3}L_{F}(x^{\prime},m;\phi),$ (7)
where $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are regularization hyperparameters that balance the contributions of the three respective terms.
In the complete loss function, we incorporate counterfactual losses $\mathcal{L}_{conEx}$ for every target class, as our goal is to enable the generation of counterfactuals in all possible directions. Additionally, we include the standard cross-entropy component $\mathcal{L}_{CE}(f(x;W),y)$ to ensure correct classification, where $y$ denotes the true label of $x$. Finally, the full loss $\mathcal{L}_{HyConEx}$ is defined as:
$\mathcal{L}_{HyConEx}(x,x^{\prime},y;\theta,\phi)=\mathcal{L}_{CE}(f(x;W),y)+\sum_{m\neq y}\mathcal{L}_{conEx}(x,x^{\prime},m;\theta,\phi).$ (8)
@@ 4.2 Initialization and training process of HyConEx
The entire training process of HyConEx consists of three phases: pre-training of the hypernetwork $H(\cdot;\theta)$, training of the conditional INF $F(\cdot;\phi)$, and fine-tuning of the hypernetwork using Equation 8. All phases are described below:
  1. 1.
    Since the hypernetwork produces not only a classification boundary but also counterfactual explanations, we need to pre-train HyConEx in a way that ensures a suitable starting point for the optimization process detailed in Section 4.1. To this end, we use clustering with the k-means algorithm [35], which partitions each class into several components. The centers of the obtained regions provide a reasonable approximation of counterfactual examples. During this phase, we utilize the classification cross-entropy loss $\mathcal{L}_{CE}(f(x),y)$, combined with the proximity loss between counterfactual candidate $x^{\prime}$ and the nearest clustering mean of the alternative class:
    $\mathcal{L}_{CE}(f(x),y)+\alpha\sum_{m\neq y}\|x^{\prime}-r_{m}\|,$
    where $x^{\prime}=x-W_{m}$ and $r_{m}$ is the closest cluster center from the $m$-th class to a data point $x$ and $\alpha$ is a hyperparameter. This clustering-based pre-training is targeted to obtain plausible counterfactuals, but with lower quality than applying INFs. However, it is more stable, which is important at the initial stage of the training.
  2. 2.
    In the second stage, we train a conditional INF $F(\cdot;\phi)$ on class labels predicted by a pre-trained hypernetwork classifier. This initialization helps to stabilize the training process and leads to faster convergence.
  3. 3.
    After pre-training stages, we freeze the parameters of INF and train only the hypernetwork network until convergence, see Algorithm 1. The inference is presented in Algorithm 2 for completeness.
Algorithm 1 HyConEx Training Iteration
Input: Hypernetwork $H(x;\theta)$ returning a linear model $f(x;W)$ by $W=H(x;\theta)$,
Conditional normalizing flow $F(\cdot|y;\phi)$ for class $y$,
Training set $\mathcal{D}$ and set of class labels $\mathcal{C}$
for each batch $(x,y)$ in $\mathcal{D}$ do
  $W\leftarrow H(x;\theta)$ $\triangleright$ Linear model
  $\hat{y}\leftarrow f(x;W)$ $\triangleright$ Current predictions for $x$
  $loss\leftarrow\mathcal{L}_{CE}(\hat{y},y)$
  for each $m$ in $\mathcal{C}$ do
   $x^{\prime}\leftarrow x-W_{m}$ $\triangleright$ Counterfactual for $x$ and class $m$
   if $m\neq y$ then
     $loss\mathrel{+}=\mathcal{L}_{conEx}(x,x^{\prime},m;\theta,\phi)$
  Update $\theta$ with respect to lossloss via gradient optimization
Algorithm 2 HyConEx Inference
Input: Hypernetwork $H(x;\theta)$ returning a linear model $f(x;W)$ by $W=H(x;\theta)$,
Factual instance $x$ and target class $t$
$W\leftarrow H(x;\theta)$ $\triangleright$ Linear model
$\hat{y}\leftarrow f(x;W)$ $\triangleright$ Current predictions for $x$
$x^{\prime}\leftarrow x-W_{t}$ $\triangleright$ Counterfactual for $x$ and class $t$
return $\hat{y},x^{\prime}$
4.3 Categorical features
To represent categorical variables numerically, we use one-hot encoding. After applying this transformation, we could theoretically assume that we are now dealing with a fully numerical problem and expect the generated counterfactual vectors to be close to one-hot vectors in the categorical positions, because of the INF component in the loss function. However, this approach unfortunately fails and leads to validity issues, as demonstrated by the PPCEF results in Table 7.
Therefore, to address this problem, we additionally introduce special smoothing on the categorical positions using softmax with a low temperature $T$. More precisely, we modify the generated vector $x^{\prime}$ at positions from $i_{1}$ to $i_{K}$, which correspond to the one-hot encoded range of a given categorical variable, as follows:
$x^{\prime}_{i_{k}}=\frac{\exp\left(\frac{x^{\prime}_{i_{k}}}{T}\right)}{\sum_{j=1}^{K}\exp\left(\frac{x^{\prime}_{i_{j}}}{T}\right)}.$
Lower values of $T$ lead to the formation of more one-hot-like vectors. In our experiments, we found that setting $T=0.01$ produced stable and satisfactory results. Therefore, we do not treat $T$ as a tunable hyperparameter but instead fixed it at this value for all experiments.
The entire process is repeated separately for each categorical feature in the dataset. Implementing this idea allows us to produce counterfactuals with near-perfect validity without compromising metrics related to density or proximity.
@@ 5 Experiments
We evaluate HyConEx in terms of classification accuracy and counterfactual explanation quality. To our knowledge, HyConEx is the first deep learning model that returns counterfactual examples by a classification model itself.
@@ 5.1 Instantiation of HyConEx
Table 2: Hyperparameters of HyConEx.
Name Description Default value
Backbone nr_blocks Number of residual blocks 4
hidden_size Hidden layer size 256
dropout_rate Dropout probability 0.25
act_func Activation function GeLU
MAF num_layers Number of autoregressive layers 8
hidden_features Hidden dimension size 16
num_blocks_per_layer Blocks per autoregressive layer 4
dropout_rate Dropout probability 0.00
act_func Activation function ReLU
Training nr_epochs Maximal number of train epochs 1500
pretraining_epochs Number of pretraining epochs 500
warm_up_epochs Warm-up epochs for $\mathcal{L}_{conEx}$ 200
components scaling
batch_size Dataloader batch size 256
learning_rate Learning rate 5e-4
optimizer Optimization method Adam
scheduler LR scheduler LambdaLR +
CosineAnnealing
$\alpha$ Initialization phase trade-off 0.8
$\alpha_{1}$ $\mathcal{L}_{CE}$ loss weight 0.8
$\alpha_{2}$ MSE loss weight 0.1
$\alpha_{3}$ $\mathcal{L}_{F}$ loss weight 0.1
For implementing HyConEx, we use TabResNet as a hypernetwork backbone and MAF as conditional INF. TabResNet uses 4 residual blocks, a hidden size of 256, and a dropout rate of 0.25. MAF is configured with 16 hidden features, 8 layers, and 4 blocks per layer.
The initialization phase uses a trade-off parameter $\alpha=0.8$ and is continued until model convergence. In the second phase, we employ the same Adam optimizer and cosine annealing scheduler as used during pre-training. For all experiments, we linearly increase the trade-off parameters $\alpha_{i}$ from 0 to $\alpha_{1}$=0.8, $\alpha_{2}$=0.1 and $\alpha_{3}$=0.1. The duration of the second phase is determined using early stopping, which aims to maximize plausibility while ensuring high model accuracy, high counterfactual validity, and low proximity.
All details are described in Table 2.
@@ 5.2 Predictive performance
To validate the predictive performance of HyConEx, we use the benchmark from [7] with an identical experimental setup. In particular, we compare HyConEx with strong ensemble methods (Random Forest and CatBoost), classical shallow classifiers (decision trees, logistic regression), as well as deep learning models equipped with interpretability tools (TabNet [25], TabResNet, IMN [26]). Our main goal is to show that introducing counterfactual functionality to a hypernetwork classifier does not deteriorate its original performance. Hence, we are particularly focused on comparing with IMN.
We use 18 tabular datasets retrieved from the OpenML repository of varied characteristics (see Table 3 for datasets IDs). We split every dataset into train, validation and test in proportion: 60-20-20. For all experiments, we perform standard preprocessing: normalizing features and handling class imbalance via downsampling. We report the Area Under the Receiver Operating Characteristic curve (AUROC), which is well-suited for imbalanced data as well.
Table 3: Comparison of predictive performance using AUROC measure. Introducing counterfactual explanations to hypernetwork classifier does not deteriorate its predictive performance (compare HyConEx with IMN).
Dataset Dataset Decision Logistic Random TabNet TabResNet CatBoost IMN HyConEx
ID Name Tree Regression Forest
3 kr-vs-kp 0.987 0.990 0.998 0.983 0.999 0.999 0.999 0.997
12 mfeat-factors 0.938 0.999 0.998 0.995 0.999 0.999 0.999 0.999
31 credit-g 0.643 0.775 0.795 0.511 0.756 0.790 0.751 0.799
54 vehicle 0.804 0.938 0.927 0.501 0.968 0.934 0.957 0.946
1067 kc1 0.620 0.802 0.801 0.789 0.808 0.800 0.805 0.809
1461 bank-marketing 0.703 0.908 0.930 0.926 0.931 0.937 0.930 0.905
1464 blood-transfusion 0.599 0.749 0.666 0.516 0.740 0.709 0.742 0.769
1468 cnae-9 0.926 0.996 0.995 0.495 0.995 0.996 0.994 0.996
1486 nomao 0.935 0.987 0.993 0.991 0.994 0.994 0.993 0.987
1489 phoneme 0.842 0.805 0.962 0.928 0.949 0.948 0.950 0.900
1590 adult 0.752 0.903 0.917 0.908 0.915 0.930 0.915 0.872
40984 segment 0.946 0.980 0.995 0.985 0.994 0.995 0.994 0.985
41142 christine 0.626 0.742 0.796 0.713 0.782 0.822 0.775 0.725
41143 jasmine 0.749 0.850 0.880 0.823 0.860 0.870 0.865 0.835
41146 sylvine 0.910 0.966 0.983 0.974 0.982 0.988 0.981 0.979
41161 riccardo 0.857 0.995 0.999 0.997 0.998 1.000 0.998 0.994
41163 dilbert 0.873 0.994 0.999 0.998 1.000 1.000 1.000 0.992
41164 fabert 0.786 0.898 0.925 0.888 0.913 0.935 0.902 0.895
Figure 4: Mean AUROC scores of comparative methods.
The results presented in Table 3 demonstrate that HyConEx maintains overall strong performance, achieving the highest AUROC on five datasets (12, 31, 1067, 1464, 1468). Additionally, Figure 4 shows that HyConEx achieves fifth best mean AUROC score. These results reveal that additional counterfactual generation capability embedded in HyConEx leads to a slight deterioration of classification performance. However, this degradation is not high.
To evaluate the significance of HyConEx’s performance against competitive methods (measured by the AUROC values presented in Table 3), we performed a formal statistical comparison. We employed the one-sided Wilcoxon signed-rank test for paired samples and adjusted the obtained p-values using the Holm-Bonferroni correction to mitigate the risk of type I errors. As shown in Table 4, none of the benchmarked methods outperform HyConEx. However, without adjusting the p-value, Random Forest, TabResNet, CatBoost and IMN would have suggested a significant performance advantage over HyConEx.
Table 4: Results of the Wilcoxon signed-rank test evaluating the HyConEx performance measured by AUROC. Adjusted p-values represent the Holm-Bonferroni correction for multiple comparisons. Bold values correspond to p-values below the significance level threshold of $0.05$.
Method statistic raw p-value adjusted p-value
Decision Tree 0.0 1.000 1.000
Logistic Regression 45.0 0.803 1.000
Random Forest 124.5 0.047 0.223
TabNet 44.5 0.935 1.000
TabResNet 122.5 0.015 0.103
CatBoost 100.0 0.049 0.223
IMN 119.5 0.021 0.125
@@ 5.3 Counterfactual explanations
Evaluation setup
We evaluated counterfactual explanations by assessing the quality, plausibility, and computational efficiency of our HyConEx method compared to established baselines such as CEM [21], CBCE [23], CEGP [20], and PPCEF [10]. Our evaluation leverages a benchmark from [10] (with an identical setup) and expands it by incorporating additional categorical datasets.
We consider multiple scenarios, including binary classification using datasets such as Moons, Law, Audit and Heloc, multiclass classification using datasets such as Blobs, Digits and Wine, and cases with categorical features using Adult [36], Credit-A [37] and Credit-G [38] datasets. For all experiments, we perform standard preprocessing: normalizing features and handling class imbalance via downsampling. For datasets with categorical features, we introduce slight Gaussian noise to labels for efficient training, while for counterfactual examples, we apply softmax activation to produce one-hot predictions and select the class with maximum value (see Section 4.3 for details).
@@
Metrics
We quantify counterfactual quality using several standard metrics from the counterfactual explanation literature [5]. Following the evaluation protocol of PPCEF [10], Coverage (Cover.) indicates the fraction of instances for which counterfactuals are generated, while Validity (Valid.) represents the percentage of counterfactuals that successfully changed the model prediction to the desired target. To assess the minimality of the changes from the original example, we report the L1 and L2 norms for continuous features and the Hamming distance for categorical features. Plausibility is measured using a probabilistic plausibility score (P.Plaus.) [10], which indicates how many counterfactuals have an estimated log density higher than the median log density of the train set. Furthermore, we provide additional plausibility metrics such as log density (LogDens) estimated using normalizing flow model and outlier detection scores such as Local Outlier Factor (LOF) [39] and Isolation Forest (IsoForest) [40]. LOF quantifies the local density deviation of a data point relative to its neighbors (higher values indicate outliers), while IsoForest measures how easily a point can be isolated in the feature space using binary trees (lower scores indicate outliers). Finally, we report the CPU computation time in seconds to generate counterfactuals to capture the efficiency of the method.
Figure 5: Summarized results from Tables 5-7 evaluating counterfactual explanations. We report only those methods which appear in all tables.
Results
@@
Our experiments demonstrate that HyConEx consistently achieves near-perfect coverage and validity across diverse settings while offering substantial computational advantages (see Tables 5-7). Results presented in Table 5 show that for binary classification tasks, our method generates counterfactual explanations with competitive proximity metrics and high plausibility, processing the Moons dataset in merely 0.003 seconds compared to 2724.99 seconds for Wachter. CEGP offers excellent L1 and L2 proximity, though this comes at the expense of lower probabilistic plausibility and density. Conversely, CBCE is very fast and generally achieves good LOF and IsoForest outlier scores, but it consistently yields the worst outcomes on distance-based metrics. All methods among CEGP, CEM, and Wachter require hours of optimization to generate counterfactual examples for all points in the test set, making them impractical for real-time applications. PPCEF specializes in maximizing P.Plaus., and its computational cost remains relatively moderate.
In the multi-class case ($K$-class scenario), we consider a setting where, for a given factual point, we create K1K-1 counterfactual points corresponding to all available target classes. Table 6 presents the resulting metrics aggregated over all classes (we omit the CEM method, as it does not support targeted counterfactuals). Noteworthy findings in the multi-class case, beyond those reported in Table 5, include the competitive performance of the Wachter method on L1 and L2 metrics, as well as the top scores achieved by HyConEx on LogDens, LOF, and IsoForest. Our main advantage—time efficiency—becomes even more important here, as the number of possible alternative classes increases (e.g., up to 9 in the case of Digits).
The comparison on datasets with categorical features is illustrated in Table 7 (we exclude the Wachter model since it is not designed for datasets with categorical features). Although both HyConEx and PPCEF use INFs to estimate data density, only our method does it correctly for categorical variables (compare validity scores and see Section 4.3 for details). HyConEx once again maintains strong performance, generating valid counterfactuals that require only minimal modifications and remain within high-density regions of the training data distribution. Among the remaining approaches, an interesting observation can be made for CEGP, which aggressively minimizes the Hamming distance and deteriorates other metrics (see low coverage for the Adult dataset).
The results aggregated over all datasets from Tables 5-7 are shown in Figure 5. Notably, HyConEx achieves very good results on in-distribution metrics (P.Plaus., LogDens, LOF, IsoForest) being significantly faster than other methods. Nevertheless, the proximity of counterfactuals generated by HyConEx is frequently lower than the ones produced by competitive methods.